Problem 13-7

Five years ago Hemingway Inc. issued 6,000 30-year bonds with par values of $1,000 at a coupon rate of 6%. The bonds are now selling to yield 5%. The company also has 15,000 shares of preferred stock outstanding that pay a dividend of $6.50 per share. These are currently selling to yield 13%. Its common stock is selling at $21, and 200,000 shares are outstanding. Assume that the coupon payments are semi-annual. Calculate Hemingway's market value based capital structure. Round the values to the nearest dollar and the weights to two decimal places. Round PVF and PVFA values in intermediate calculations to four decimal places. Do not round other intermediate calculations.

Component Value Capital Structure
Debt $   %
Preferred stock %
Equity %
Total Capital $   %
Answer #1
                  K = Nx2
Bond Price =∑ [(Semi Annual Coupon)/(1 + YTM/2)^k]     +   Par value/(1 + YTM/2)^Nx2
                   k=1
                  K =25x2
Bond Price =∑ [(6*1000/200)/(1 + 5/200)^k]     +   1000/(1 + 5/200)^25x2
                   k=1
Bond Price = 1141.81

Preferred equity price = dividend/yield = 6.5/0.13 = 50

MV of equity=Price of equity*number of shares outstanding
MV of equity=21*200000
=4200000
MV of Bond=Par value*bonds outstanding*%age of par
MV of Bond=1000*6000*1.14181
=6850860
MV of Preferred equity=Price*number of shares outstanding
MV of Preferred equity=50*15000
=750000
MV of firm = MV of Equity + MV of Bond+ MV of Preferred equity
=4200000+6850860+750000
=11800860
Weight of equity = MV of Equity/MV of firm
Weight of equity = 4200000/11800860
W(E)=0.3559 = 35.59%
Weight of debt = MV of Bond/MV of firm
Weight of debt = 6850860/11800860
W(D)=0.5805 = 58.05%
Weight of preferred equity = MV of preferred equity/MV of firm
Weight of preferred equity = 750000/11800860
W(PE)=0.0636 = 6.36%
Component Value Capital Structure
Debt 6850860 58.08
Preferred stock 750000 6.36
Equity 4200000 35.59
Total Capital 11800860 100